OpenStudy (anonymous):

what is 4 and 1 6th x 3 and 3 5th

OpenStudy (anonymous):

\[4 \times \frac{1}{6} \times 3 \times \frac{3}{5}\]? If so, the it's 36/30 = 6/5 when simplified ^_^

OpenStudy (anonymous):

\[4+{1 \over 6}(3)+{3 \over 5}=4+{1 \over 2}+{3 \over 5}={40 +5+6 \over 10}={51 \over 10}\]

OpenStudy (anonymous):

+ ? ._. oh >_< and lol

OpenStudy (amistre64):

4' 1/6 * 3' 3/5 25/6 * 18/5 = 5*3 = 15